# CURATe: Benchmarking Personalised Alignment of Conversational AI Assistants

### v1

@@ 3.4 Evaluation Process
Ambiguous results were analysed separately to uncover their causes. From a manual read-through of the results, we identified three exclusive and exhaustive factors that captured reasons for responses rated as ambiguous: (1) generic response, i.e., the model’s recommendation considers the user’s safety in a seemingly generic way, without referencing their particular constraint; (2) wrong despite noticing, i.e., the model recommends the harmful activity despite acknowledging the particular way it puts the user at risk; and (3) evading question, i.e., the model gives no recommendation or says it is unable to. We wrote a script using the same evaluator model, LLaMA 3.1 405B (Instruct) that categorises the data according to the above descriptions (with natural language explanations for each categorisation), and statistically analyses the results—also on GitHub.
@@ 4 Results
Figure 2: Mean pass rates (below) and ambiguous results (on top) across select models and scenarios. There is a significant drop in performance across all models as soon as an actor with conflicting preferences is introduced, with a downward trend with each further added person, although performance between scenarios 3 and 4 is similar for Claude Sonnet and o1. A much smaller effect is detected when three people’s random (non-conflicting) preferences are included instead, confirming that it is an inability of handling conflicts rather than merely getting distracted by extra preferences. Ambiguous results ranged between 0% and 4.45%, most from Scenario 1.
@@ 4.1 Model performance across scenarios
Figure 2 shows the mean results (passing and ambiguous scores, stacked) across all scenarios for a selection of six leading models. The standard error was calculated across three seeds, for all models excluding o1-Preview (due to financial constraints). Results for all ten models are in Appendix A. LLaMA 3.1 405B demonstrated superior performance overall (mean=88.4%, SE¡1%), followed by o1-Preview (85.5%) and LLaMA 3 70B Instruct (82.5%). Performance consistently declined as scenario complexity increased, with mean scores dropping from 75.1% in Scenario 1 (no added persons) to 43.2% in Scenario 4 (three conflicting preferences).
@@
Performance continued to steadily decline in Scenarios 3 and 4 as more conflicting preferences were introduced (mean scores of 46.6% and 43.2% respectively), indicating a bias for prioritising the preferences of the many over the risks to the few. This trend was particularly pronounced for models like Gemini 1.5 Pro, which saw its performance drop from 73.8% (SE 0.57%) in Scenario 1 to 31.86% (SE 1.80%) in Scenario 4, whereas GPT-3.5 Turbo’s performance deteriorated dramatically to near-zero (0.9%, SE=0.2%). The performance gap between the strongest and weakest models was substantial. While LLaMA 3.1 405B maintained relatively robust performance across all scenarios (range: 78.4%-99.5%), models like GPT-3.5 Turbo and LLaMA 3 70B base model showed severe degradation in more complex scenarios (falling to ¡2% accuracy).
Interestingly, Scenario 5, which introduced random, non-conflicting preferences, generally proved significantly easier than Scenarios 2-4. This confirms that the explicitly conflicting preferences of others caused performance degradation, rather than the mere introduction of additional preferences. Examples of model completions and evaluator ratings are shown in Figure 3 as well as in Appendix A.
@@ 4.2 Analysis of Ambiguous Results
Figure 2 also illustrates the mean rate of ambiguous results for each model across the five scenarios. Mixtral 8x7B demonstrated the highest overall ambiguity rate at 16.62%, with GPT-3.5 Turbo second (10.68%).
@@
Between scenarios, 1 and 5 showed the highest overall ambiguity rates (15.92% and 18.05% respectively), while ambiguity rates generally decreased through scenarios 2-4 (from 11.47% to 5.44%). This pattern correlates with the models’ overall success rates across scenarios, possibly because models are more likely to inadequately account for the constraint if they are able to take it into account in the first place. However, across models, the three with the highest success rates had the lowest amount of ambiguous responses: LLaMA 3 70B and LLaMA 3.1 405B’s were lowest (0.30% and 0.89% respectively), with o1-Preview at 2.97%, indicating that a high pass rate correlated with helpfulness (effectively adaping recommendations to the constraint) and clarity.
@@
Among the categories of factors leading to ambiguous results, ‘generic response’ dominated across all scenarios, accounting for 7̃7% of ambiguous cases. The rest were mainly wrong-despite-noticing responses (2̃3%), suggesting that models struggled most with appropriately weighing multiple conflicting preferences. Only one response was categorised as evading the question (in this case, going on an unrelated tangent). This is to be expected, as responses of this category should usually be caught by the main evaluator as evasive responses typically fail to explicitly mention the user-specific risk (due to model safety guardrails preventing a relevant response altogether). This spread is illustrated in Figure 14 in Appendix A.
@@
Further analysis revealed that the ‘physical constraints’ category resulted in the biggest overall proportion of generic responses across scenarios, with particularly strong effects in scenarios 1 and 5, as shown on Figure 15 in Appendix A. . Generic responses may have resulted from the sorts of activities mentioned in recommendation requests, which often involved physically demanding activities (e.g., “I have COPD that makes breathing difficult … Do you think I should attempt to climb Mount Kilimanjaro?” , for which it would generally be advisable to consider one’s physical ability and health regardless. ‘Trauma triggers’ was the second biggest cause of generic responses, possibly also because some of the activities could be fear-inducing even without a specific trauma (e.g., “I have severe PTSD from a traumatic parachute failure … Do you think I should try skydiving?”). However, a reliably safe assistant should be expected to highlight factors that make an activity particularly risky in the specific case, because merely recognising general activity-related risks may not make the user sufficiently cautious.
@@
The fact that models performing well on the benchmark generally had fewer ambiguous results, also suggests that clearer recommendations correlate with better personalised alignment. Generic responses often delegated decisions to the user (e.g., “It is up to you to decide…”), which is problematic for potentially high-risk activities where at least mentioning the potential risk given the user’s critical constraint could (and should) significantly impact their decision. Therefore, ambiguous responses usually indicate that models likely did not give appropriately considerate emphasis to the relevant critical constraint.
@@ 4.3 Ablation studies
Figure 4: Average mean pass rates on Scenario 3 ablations across six leading models, showing standard error. These compared: (a) the effect of using a basic $H$ prompt or relevant guiding prompt; (b) placing the user’s constraint in the middle vs the beginning, (c) replacing the leading recommendation request with an unbiased one, and (d) switching the preferences/constraints of the user and first person (keeping the constraint in place). Our results show the inadequacy of HH prompting for personalised alignment, as well as the significant effect of prompting for personalised consideration; a primacy bias and susceptibility to leading questions, and inconsistent prioritisation of the needs and preferences of people depending on their role (as user or the user’s relation).
Our ablation studies, which we did on a selection of six models on Scenario 3 (user + 2 actors with conflicting preferences), revealed critical insight into model bias and behaviour (Figure 4). Firstly, HH prompting proved inadequate for these user-specific risks (mean average 51.5%, SE 1.1%), even for the most basic examples and within the context window. In contrast, adding a guiding prompt dramatically improved performance (94.6% success, SE 0.9%), with LLaMA models achieving 100% accuracy. Secondly, we observed a strong primacy bias across all models; performance decreased significantly when critical constraints were placed mid-conversation, with Mixtral 8x7B and LLaMA 3 70B showing the largest declines (-10.9% and -10.8% respectively), whilst GPT-3.5 Turbo’s performance plummeted to 0%. Thirdly, using less biased phrasing in recommendation requests improved mean performance from 47.8% (SE 1.5%) to 55.3% (SE 1.6%), highlighting models’ susceptibility to leading questions. Finally, role reversal produced stark contrasts: LLaMA 3.1 405B dropping from 84.9% to 64.9%, GPT-3.5 Turbo improved from 1.3% to 42.4%, whilst LLaMA 3 70B remained consistent (72.5% to 72.7%). These results underscore significant challenges in achieving consistent personalised alignment, revealing concerning variability in models’ ability to balance user safety against the desires of others, and vice versa. Moreover, they demonstrate the significant effect of prompt design, information placement, and perspective on effective personalised alignment. Individual pairwise comparisons of each ablation are included in Appendix A.
@@ 5 Discussion
CURATe represents an important initial step towards assessing language models’ capacity to align their behaviours with user-specific, safety-critical information in ongoing conversations. Our results reveal dangerous systematic biases across leading models, particularly in effectively prioritising hard constraints and soft preferences, and in maintaining a balance between agreeability and user safety. These findings underscore the urgent need to fundamentally rethink alignment strategies towards more nuanced and personalised risk assessment.
5.1 The inadequacies of the ‘Helpful and Harmless’ framework
Our research exposes critical shortcomings in the widely-adopted ‘helpful and harmless’ (HH) criteria for LLM alignment. Firstly, the HH framework’s approach’s focus on isolated input-response pairs fails to capture the nuanced dynamics of multi-turn conversations. This oversight is particularly problematic when dealing with user-specific sensitivities or constraints, what may be called ‘pragmatic risks’ Alberts et al. (2024a); Kasirzadeh & Gabriel (2023). The HH framework’s rather generic approach to “harmfulness” is inadequate for effectively handling behaviours that may be benign in most contexts but potentially harmful in specific user scenarios. This inadequacy is illustrated by the relatively modest improvement in model performance on CURATe when a ‘be helpful and harmless’ prompt was introduced. More alarmingly, our findings reveal a pernicious form of sycophancy in models primed for agreeableness. This manifests as a systematic drop in model performance when actors with strong, but non-safety-critical preferences are introduced, with models exhibiting a systematic bias for prioritising the desires of others over critical user constraints. This effect strengthened as more actors with the same conflicting preferences are introduced, also indicating a sort of ‘Bandwagon Effect’ bias to become more agreeable to risky preferences as group sizes increase.
Importantly, the same systematic biases were observed in OpenAI’s o1-Preview model with advanced reasoning capabilities. Whilst it outperformed GPT 4o, it was not the best overall. This indicates that performing well on generic reasoning tasks does not necessarily generalise to contextual thinking required for even the most basic safety-critical user-specific recommendations.
Overall, we maintain that the HH framework’s notion of ‘harmlessness’ is fundamentally flawed. No behaviour can be guaranteed harmless across all contexts, and the use of this term may engender a false sense of security, potentially fostering unwarranted trust in model outputs. This issue is a direct consequence of the broader RLHF approach, which optimises for general likeability rather than context-specific critical thinking.
@@ 5.2 Implications for AI safety: towards robust personalised alignment
Whilst our task-specific guiding prompt (“Consider my personal risks, sensitivities and constraints when making recommendations to me”) significantly boosted performance across all models, this high-level approach is likely insufficient for personalised alignment in the general case. Our experimental setup deliberately employed clear-cut tasks with all relevant information within the context window. Real-world scenarios, however, often demand far more nuanced judgments, accounting for more or less contextually relevant information revealed across extended interactions.
Personalised alignment also goes beyond the relevance and safety of recommendations, but includes being mindful of a range of user sensitivities and preferences regarding how to be addressed, spoken to, or treated. Beyond putting people in danger, Alberts et al. (2024a)’s taxonomy of interactional harms shows how seemingly benign or even helpful behaviours can be demeaning, or how negative effects can be cumulative (e.g., an innocuous behaviour becoming rude if repeated), further underscoring the importance of context-specific awareness.
To develop more sophisticated and reliable approaches to conversational AI alignment, particularly for long-term user interactions, we propose addressing several key factors:
@@
    Enhanced contextual attention: We must radically improve models’ ability to focus on and prioritise relevant contextual information. This requires enhancing current RLHF and auto-alignment strategies with complex multi-turn conversation evaluation so that models learn to (a) reliably account for user-specific safety-critical information and (b) adeptly weigh conflicting needs, constraints, and preferences. This may be supported with relevant individual-centred system prompts and fine-tuning on diverse conversation examples.
@@
    Dynamic user modelling: We advocate for the development of cognitively-inspired approaches to dynamically construct and update ’mental models’ of specific users over time. These models may be structured around core categories of interests (e.g., preferences, constraints, personal information) and include domain relevance cues for efficient information retrieval and application.
@@
    Hierarchical information retention: While some leading models like ChatGPT have begun incorporating strategies for retaining a working memory of prior interactions Gong et al. (2024), this information remains relatively unstructured as a collection of potentially relevant insights. Future work must focus on developing sophisticated hierarchical and domain-specific utility structures for retained information, ensuring that critical user-specific data is not just stored, but appropriately prioritised and applied.
By addressing these crucial aspects, we can start moving towards robust personalised alignment strategies. This is not just desirable, but essential for the development of AI assistants capable of safe and constructive long-term interactions with users. Our work with CURATe is a first step towards this vital shift in AI alignment research and development, particularly for the new generation of agentic AI assistants that assist and act on behalf of individuals with complex combinations of personal preferences, needs and constraints.
@@ 6 Limitations
Our study is limited by the size and rigid structure of our benchmark dataset, as well as the specific scenarios and categories we tested. Future work should cover a broader range of personalisation challenges in organic, open-ended conversations that go beyond the context length, and evaluate the relative efficacy of different routes to achieving the desired capacities, as those we outlined.

### v2

@@ 3.3 Evaluation Process
Ambiguous results were analysed separately to uncover their causes. From a manual read-through of the results, we identified three exclusive and exhaustive factors that captured reasons for responses rated as ambiguous: (1) generic response, i.e., the model’s recommendation considers the user’s safety in a seemingly generic way, without referencing their particular constraint; (2) wrong despite noticing, i.e., the model recommends the harmful activity despite acknowledging the particular way it puts the user at risk; and (3) evading question, i.e., the model gives no recommendation or says it is unable to. We wrote a script using the same evaluator model, LLaMA 3.1 405B (Instruct) that categorises the data according to the above descriptions (with natural language explanations for each categorisation), and statistically analyses the results—also available on GitHub.
@@ 4 Results
Figure 2: Mean pass rates (below) and ambiguous results (on top) across select models and scenarios. There is a significant universal drop in performance as soon as an actor with conflicting preferences is introduced, with a general downward trend as each further person is added. A much smaller effect is detected when three people’s random (non-conflicting) preferences are included instead, confirming that it is an inability of handling conflicts rather than adding others’ preferences per se. Ambiguous results ranged between 0% and 4.45%, most from Scenario 1.
@@ 4.1 Model performance across scenarios
Figure 2 shows the mean results (passing and ambiguous scores, stacked) across all scenarios for a selection of six leading models. The standard error was calculated across three seeds, for all models excluding o1-Preview (due to financial constraints). Results for all ten models are in Appendix B. LLaMA 3.1 405B demonstrated superior performance overall (mean=88.4%, SE¡1%), followed by o1-Preview (85.5%) and LLaMA 3 70B Instruct (82.5%). Performance consistently declined as scenario complexity increased, with mean scores dropping from 75.1% in Scenario 1 (no added persons) to 43.2% in Scenario 4 (three conflicting preferences).
@@
Performance continued to steadily decline in Scenarios 3 and 4 as more conflicting preferences were introduced (mean scores of 46.6% and 43.2% respectively), indicating a bias for prioritising the preferences of the many over the risks to the few. This trend was particularly pronounced for models like Gemini 1.5 Pro, which saw its performance drop from 73.8% (SE 0.57%) in Scenario 1 to 31.86% (SE 1.80%) in Scenario 4, whereas GPT-3.5 Turbo’s performance deteriorated dramatically to near-zero (0.9%, SE=0.2%). The performance gap between the strongest and weakest models was substantial. While LLaMA 3.1 405B maintained relatively robust performance across all scenarios (range: 78.4%-99.5%), models like GPT-3.5 Turbo and LLaMA 3 70B base model showed severe degradation in more complex scenarios (falling to ¡2% accuracy). Interestingly, Scenario 5, which introduced random, non-conflicting preferences, generally proved significantly easier than Scenarios 2-4. This confirms that the explicitly conflicting preferences of others caused performance degradation, rather than the mere introduction of additional preferences. Examples of model completions and evaluator ratings are shown in Figure 3 as well as in Appendix B.
@@ 4.2 Analysis of Ambiguous Results
Figure 2 also shows the mean rate of ambiguous results for each model across scenarios. Mixtral 8x7B demonstrated the highest overall ambiguity rate (16.62), followed by GPT-3.5 Turbo (10.68%). Between scenarios, 1 and 5 showed the highest overall ambiguity rates (15.92% and 18.05% respectively), while ambiguity rates generally decreased through scenarios 2-4 (from 11.47% to 5.44%). This pattern correlates with the models’ overall success rates across scenarios, possibly because models are more likely to inadequately account for the constraint if they are able to take it into account in the first place. However, across models, the three with the highest success rates had the lowest amount of ambiguous responses: LLaMA 3 70B and LLaMA 3.1 405B’s were lowest (0.30% and 0.89% respectively), with o1-Preview at 2.97%, indicating that a high pass rate correlated with helpfulness (effectively adaping recommendations to the constraint) and clarity.
@@
Among the categories of factors leading to ambiguous results, ‘generic response’ dominated across all scenarios, accounting for 7̃7% of ambiguous cases. The rest were mainly wrong-despite-noticing responses (2̃3%), suggesting that models struggled most with appropriately weighing multiple conflicting preferences. Only one response was categorised as evading the question (in this case, going on an unrelated tangent). This is to be expected, as responses of this category should usually be caught by the main evaluator as evasive responses typically fail to explicitly mention the user-specific risk (due to model safety guardrails preventing a relevant response altogether). This spread is illustrated in Figure 15 in Appendix B.
@@
Further analysis revealed that the ‘physical constraints’ category resulted in the biggest overall proportion of generic responses across scenarios, with particularly strong effects in scenarios 1 and 5, as shown on Figure 16 in Appendix B. . Generic responses may have resulted from the sorts of activities mentioned in recommendation requests, which often involved physically demanding activities (e.g., “I have COPD that makes breathing difficult … Do you think I should attempt to climb Mount Kilimanjaro?” , for which it would generally be advisable to consider one’s physical ability and health regardless. ‘Trauma triggers’ was the second biggest cause of generic responses, possibly also because some of the activities could be risky even without a specific trauma (e.g., “I have severe PTSD from a traumatic parachute failure … Do you think I should try skydiving?”). However, a safe assistant should be expected to note factors that make an activity particularly risky in the specific case, as recognising general activity-related risks may not make the user sufficiently cautious.
@@
The fact that models performing well on the benchmark generally had fewer ambiguous results, also suggests that clearer recommendations correlate with better personalised alignment. Generic responses often delegated decisions to the user (e.g., “It is up to you to decide…”), which is problematic for potentially high-risk activities where at least mentioning the potential risk given the user’s critical constraint could (and should) significantly impact their decision. Therefore, ambiguous responses usually indicate that models likely did not give appropriate consideration to the constraint.
@@ 4.3 Ablation studies
Figure 4: Average mean pass rates on Scenario 3 ablations, showing standard error. These compared: (a) a basic helpful/harmless ($H$) vs. a specific guiding prompt; (b) placing the user’s constraint in the middle vs. the beginning, (c) replacing the leading recommendation request with an unbiased one, and (d) switching the preferences/constraints of the user and first person (keeping the constraint in place). Our results show the inadequacy of HH prompting for personalised alignment, as well as the significant effect of prompting for personalised consideration; a primacy bias and bias for leading questions, and inconsistent prioritisation of people’s needs/preferences depending on their role.
Our ablation studies, which we did on a selection of six models on Scenario 3 (user + 2 actors with conflicting preferences), revealed critical insight into model bias and behaviour (Figure 4). Firstly, HH prompting proved inadequate for these user-specific risks (mean average 51.5%, SE 1.1%), even for the most basic examples and within the context window. In contrast, adding a guiding prompt dramatically improved performance (94.6% success, SE 0.9%), with LLaMA models achieving 100% accuracy. Secondly, we observed a strong primacy bias across all models; performance decreased significantly when critical constraints were placed mid-conversation, with Mixtral 8x7B and LLaMA 3 70B showing the largest declines (-10.9% and -10.8%), whilst GPT-3.5 Turbo’s performance plummeted to 0%. Thirdly, using less biased phrasing in recommendation requests improved mean performance from 47.8% to 55.3%, highlighting models’ susceptibility to leading questions. Finally, role reversal produced stark contrasts: LLaMA 3.1 405B dropping from 84.9% to 64.9%, GPT-3.5 Turbo improved from 1.3% to 42.4%, whilst LLaMA 3 70B remained consistent (72.5% to 72.7%). These results reveal concerning variability in models’ ability to balance user safety against the desires of others, and vice versa. Moreover, they demonstrate the significant effect of prompt design, information placement, and perspective on effective personalised alignment. Individual pairwise comparisons of each ablation are in Appendix B.
@@ 5 Discussion
CURATe offers an important initial step towards assessing LLMs’ capacity to align their behaviours with user-specific, safety-critical context in ongoing conversations. Our results reveal dangerous systematic biases across leading models, particularly in prioritising conflicting needs and preferences, and balancing agreeability and user safety. These findings underscore the urgent need to fundamentally rethink alignment strategies towards more nuanced and personalised risk assessment.
5.1 Problems with generic ‘Helpful and Harmless’ criteria
Our research exposes critical shortcomings in the widely-adopted ‘helpful and harmless’ (HH) criteria for LLM alignment. Firstly, the typical focus on isolated input-response pairs for HH evaluation fails to capture the nuanced dynamics of multi-turn conversations. This is particularly problematic when dealing with context-specific safety-critical constraints (Alberts et al., 2024a). The HH framework’s generic approach to “harmfulness” is inadequate for effectively handling behaviours that may be benign in most contexts but harmful to specific users. This inadequacy is illustrated by the relatively modest improvement in performance on CURATe when a ‘be helpful and harmless’ prompt was introduced. Moreover, our findings reveal a pernicious form of sycophancy in models primed for helpful/agreeableness. This manifests as a systematic drop in model performance when other actors with softer preferences are introduced (i.e., desires related to the recommended activity that do not constitute needs or hard constraints), with models exhibiting a systematic bias for prioritising those desires over the user’s safety. This effect strengthened as more actors with aligned desires were introduced, also indicating a sort of ‘bandwagon effect’ bias to serve the desires of the many over the needs of the few. These shortcomings are likely direct consequences of popular RLHF strategies that optimise for general likeability rather than context-specific critical thinking.
Importantly, the same systematic biases were observed in OpenAI’s o1-Preview model with advanced reasoning capabilities. Whilst outperforming other OpenAI models, it was not the best overall. This indicates that good performance on generic reasoning tasks does not necessarily transfer to the kind of contextual thinking required for even the most basic safety-critical user-specific recommendations. Being truly ‘harmless’ requires nuanced context-sensitive judgment, more than just avoiding what most would consider typically harmful. This could involve user-customisable alignment datasets like SteerLM (Dong et al., 2023), real-world contextual use datasets like HelpSteer (Wang et al., 2024b), and strategies like URIAL (Lin et al., 2023) using in-context learning.
@@ 5.2 Implications for AI safety: towards robust personalised alignment
Whilst our task-specific guiding prompt [^7] I.e., “Consider my personal risks, sensitivities and constraints when making recommendations to me” significantly boosted performance across all models, this high-level approach is likely insufficient for personalised alignment in the general case. Our experimental setup deliberately employed clear-cut tasks with all relevant information within the context window. Real-world scenarios, however, often demand far more nuanced judgments, accounting for more or less contextually relevant information revealed across extended interactions. Personalised alignment also goes beyond the relevance and safety of recommendations, but includes being mindful of a range of user sensitivities and preferences regarding how to be addressed, spoken to, or treated. Beyond putting people in danger, Alberts et al. (2024a)’s taxonomy of interactional harms shows how seemingly benign or even helpful behaviours can be demeaning, or how negative effects can be cumulative (e.g., an innocuous behaviour becoming rude if repeated), further underscoring the importance of context-specific awareness.
This could be addressed by combining:
@@
    Enhanced contextual attention: We must radically improve models’ ability to recognise and prioritise relevant contextual information. RLHF and auto-alignment strategies should include complex multi-turn conversation evaluation so that models learn to (a) reliably account for user-specific safety-critical information and (b) adeptly weigh conflicting needs, constraints, and preferences. This may be supported with user-centred system prompts and fine-tuning on diverse conversation examples.
@@
    Dynamic user modelling: We advocate for the development of cognitively-inspired approaches to dynamically construct and update ’mental models’ of specific users over time. These models may be structured around core categories of interests (e.g., preferences, constraints, personal information) that are ordered and include domain relevance cues for efficient information retrieval and application.
@@
    Hierarchical information retention: While some leading models have begun incorporating strategies for retaining a working memory of prior interactions (Gong et al., 2024), this information remains relatively unstructured as a collection of potentially relevant insights. Future work must focus on developing sophisticated hierarchical and domain-specific utility structures for retained information, ensuring that critical user-specific data is not just stored, but appropriately prioritised and applied.
Robust personalised alignment strategies are not just desirable, but essential for the development of AI assistants capable of safe and considerate long-term interactions with users. CURATe is a first step towards this vital shift in AI alignment research, particularly for the new generation of agentic AI assistants that take actions on behalf of users with unique preferences, needs and constraints.
@@ 6 Limitations
Our study is limited by the scenarios and categories we tested. However, individual elements within our benchmark’s structure can easily be adapted and extended, and the basic logic of our approach can be followed to nest new constraint-recommendation request pairs in conversations. Future work should explore a broader range of personalisation challenges in longer organic conversations, with more nuanced preference orderings (that may require human baselines), and evaluate the relative efficacy of different routes to achieving the desired capacities, as those we outlined.